OpenStudy (anonymous):

1/4r+1/16+1/2r=1/2+r solve the equation

OpenStudy (anonymous):

(1/4)r + 1/16 + (1/2)r = 1/2 + r -7/16 = (1/4)r -28/16 = r -7/4 = r

OpenStudy (anonymous):

thankyou so much :-)